Packages of sugar bags for Sweeter Sugar Inc. have an average weight of 16 ounces and a standard deviation of 0.2 ounces

9. Packages of sugar bags for Sweeter Sugar Inc. have an average weight of 16 ounces and a standard deviation of 0.2 ounces. The weights of the sugar packages are normally distributed. What is the probability that 16 randomly selected packages will have a weight in excess of 16.075 ounces? 

A. 0.9332 B. 0.9110 C. 0.3520 D. 0.0668 E. 0.0500 


10. Suppose that 50 percent of the voters in a particular region support a candidate. Find the probability that a sample of 1,000 voters would yield a sample proportion in favor of the candidate within 4 percentage points of the actual proportion. 

A. 0.9886 B. 0.9600 C, 0.9484 D. 0.0155 E. 0.0098
Answer #1

Ans:

9)

z=(16.075-16)/(0.2/sqrt(16)))

z=1.5

P(z>1.5)=0.9332

10)

z=0.04/SQRT(0.5*(1-0.5)/1000)=2.53

P(-2.53

=0.9943-0.0057=0.9886

Option A is correct(0.9886)